OpenStudy (anonymous):

Help Please!! Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2 y - 4 = 1/2 (x - 1) y + 4 = 2(x - 1) y + 4 = 1/2 (x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

Well, it looks like all your answers are in point slope form. The equation is : y-y1=m(x-x1) Can you figure it out, now? :o)

OpenStudy (anonymous):

@Butterfly16 its just im really confused on how to do it i really appreacite u trying to help me but im just really confused in it. i think it answer #2. im i right?

OpenStudy (anonymous):

Well, m=slope, and x1 and y1 are the points you're given. So you just simply plug them in: y-(-4)=1/2(x-1) y+4=1/2(x-1) Does that make sense?

OpenStudy (anonymous):

@Butterfly16 kinda... So i think its number 1 now? im i right?

OpenStudy (anonymous):

I gave you the answer, you just have to match it. Is there any part you dont understand?

OpenStudy (anonymous):

@Butterfly16 ohhhhh i get it now!! Thank you!!!